OpenStudy (anonymous):

Find a power series representation for f(x) = ln((1+x)/(1-x)). f(x) = sum n=0 to infinity of ?

OpenStudy (amistre64):

might want to split the log into 2 parts

OpenStudy (amistre64):

ln(a/b) = ln(a) - ln(b) now taking derivatives if need be is suitable

OpenStudy (anonymous):

ln(1+x)-ln(1-x) d/dx = (1/1+x)-(1/1-x)

OpenStudy (amistre64):

a shortcut now is to do the long division to come up with a series for those derivatives

OpenStudy (amistre64):

OR, combine them in that state and do the division

OpenStudy (anonymous):

oops (-1/1-x)

OpenStudy (anonymous):

ok, ill try.

OpenStudy (amistre64):

\[\frac{(1-x)+(1+x)}{1-x^2}\] \[\frac{2}{1-x^2}\]

OpenStudy (anonymous):

I dont understand what im supposed to be doing when you said 'do the division'

OpenStudy (amistre64):

umm, divide :/ \[\frac{2}{1-x^2}=2(\frac{1}{1-x^2})\] so 1+x^2+x^4+x^6+..... ----------------- 1-x^2 | 1 -(1-x^2) --------- x^2 -(x^2-x^4) --------- x^4

OpenStudy (amistre64):

so the derivative of the stated problem is:\[2(1+x^2+x^4+x^6+...)\] therefore its integration should be equal to the stated problem ... \[2(\frac11x^1+\frac13x^3+\frac15x^5+\frac17x^7+...)\]

OpenStudy (amistre64):

do we need to go over how to do long division?

OpenStudy (anonymous):

No, I know how. I was just confused as to why/what to divide.

OpenStudy (anonymous):

I don't believe my professor showed us how to solve these problems like this.

OpenStudy (amistre64):

in alot of the cases in a textbook, division will produce an equivalent function in polynomial form

OpenStudy (amistre64):

for example\[\frac19\] lets say 9 = 1-x \[\frac{1}{1-x}\] we are still doing the division of: --------- 9 | 1 or written another way --------- 1-x | 1 1+x+x^2+x^3+x^4+.... --------- 1-x | 1 -1(1-x) ------- x -x(1-x) ------- x^2 -x^2(1-x) --------- x^3

OpenStudy (amistre64):

9=1-x when x=-8 therefore: \(1-8+8^2-8^3+8^4_...=\frac19\)

OpenStudy (anonymous):

Oh, ok. I'm following

OpenStudy (amistre64):

thats all i got :)

OpenStudy (amistre64):

kids want food ... so bye

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

When you did the derivative, you're in an okay position to start finding the power series representations. \[\begin{align*}\frac{d}{dx}f(x)&=\frac{1}{1+x}-\frac{1}{1-x}\\&=\frac{1}{1-(-x)}-\frac{1}{1-x}\\ &=\sum_{n=0}^\infty (-x)^n-\sum_{n=0}^\infty x^n \end{align*}\] From here, you could simplify by seeing which terms if any cancel out.